the mean age of a and b five years ago was 15 years . mean age og a,b and c today is 20 years . how old will c be after 10 years ?

Answer by mananth(16949)

the mean age of a and b five years ago was 15 years . mean age og a,b and c today is 20 years . how old will c be after 10 years ?
let the ages of a & b be x &y respectively.
x-5, y-5 will be their ages 5 years ago.
Their average = 15
x-5+y-5 /2 =15
x+y-10 / 2 = 15
x+y=30+10
x+y=40
.
Let the age of c be z years now
Their average is 20
x+y+z / 3 = 20
Plug the value of x+y
40+z = 60
z=20
z will be 20+10 = 30 years after 10 years.